A student standing on a train station observes that the pitch of the train decreases from 900.3 Hz to 880.3 Hz as it travels past the student at a constant speed. What's the speed of the train? (Assume the velocity of sound is 333 m/s).
Answer #1

The question is based on the Dopplers effect. The relation between the apparent t and the actual frequency when the train is
Answer #2

Given that

frequency f=900.3 Hz

appreant frequency f'=880.3 Hz

now we find the speed of the train

f/f'=v+vs/v-vs

900.3/880.3=333+Vs/333-Vs

299799.9-900.3Vs=293139.9+880.3vs

6660=1780.3Vs

Vs=3.741 m/s